# Quantum density estimation with density matrices: Application to quantum anomaly detection

### 22 Dec 2023 v4

@@ I Introduction
In this article, we present a novel non-parametric quantum-classical density matrix density estimation strategy, called Q-DEMDE, suitable for current noisy quantum computers. It combines a proposed quantum algorithm for computing the expectation values of density matrices with a novel quantum representation of data called quantum Fourier features. The proposed method, unlike previous quantum density estimation methods [19, 20], builds probability distributions of data in the form of mixed quantum states in qubit-based quantum hardware. It is based on the density matrix kernel density estimation (DMKDE) algorithm [27] originally developed for classical computers; it fuses density matrices with random Fourier features (RFF) [28] to build a nonparametric density estimator. We show that the proposed Q-DEMDE method can be implemented in current IBM quantum computers.
@@ II Related Work
Density estimation is an important challenge in both classical and quantum machine learning. Both parametric and non-parametric methods for DE have been widely explored in classical computers [14, 15, 16, 11, 12, 13]. Some early works on quantum density estimation build parametric models based on multivariate Gaussian distributions on quantum devices [17, 18]. On the non-parametric side, the DMKDE [27] algorithm was originally developed for classical computers, but it can be implemented in quantum computers with both pure-state and mixed-state approaches; the pure-state approach has been developed in qubit-based quantum computers [20], while the mixed-state version has been implemented in a high-dimensional quantum computer [19]. In contrast to these previous works, we present a QDE strategy built on mixed states for present-day qubit-based quantum devices and propose a quantum algorithm for estimating the expected value of a density matrix that extend the mixed-state DMKDE [19] to qubit-based quantum hardware.
@@ III.1 Kernel Density Estimation
$p_{\gamma}(\mathbf{x^{*}})=\frac{1}{NM_{\gamma}}\sum_{j=0}^{N-1}k(\mathbf{x}_{j},\mathbf{x^{*}})=\frac{1}{NM_{\gamma}}\sum_{j=0}^{N-1}e^{\gamma\norm{\mathbf{x}_{j}-\mathbf{x^{*}}}^{2}},$ (1)
where $M_{\gamma}=(\pi/\gamma)^{-\frac{D}{2}}$ is a normalization constant. This method scales with the size of the training data set and requires storing the whole data set in memory.
@@ III.2 Random and Adaptive Fourier features
by sampling i.i.d the weights $\{\mathbf{w}_{i}\}\in\mathbb{R}^{D}$ from $\mathcal{N}(\mathbf{0},\mathbf{I})$, and i.i.d. biases $\{b_{i}\}\in\mathbb{R}$ from $\text{Uniform}[0,2\pi]$, leading to $\mathbf{z}_{\gamma}(\mathbf{x})^{T}\mathbf{z}_{\gamma}(\mathbf{y})\approx e^{-\gamma{\norm{\mathbf{x}-\mathbf{y}}}}$. This method follows from Bochner theorem [52], which states that a shift-invariant kernel is the Fourier transform of a probability measure, e.g., the Gaussian kernel with parameter $\gamma$ is the Fourier transform of a normal distribution with variance $2\gamma$. In addition, the kernel approximation can be improved by means of adaptive Fourier features [32], which learns the optimal RFF weights and biases $\{(\mathbf{w}_{i},b_{i})\}$ by minimizing the loss function, $\mathcal{L}_{\{(\mathbf{w}_{i},b_{i})\}}=\sum_{l,m}(k(\mathbf{x}_{l},\mathbf{x}_{m})-\mathbf{z}_{\gamma}(\mathbf{x_{l}})^{T}\mathbf{z}_{\gamma}(\mathbf{x}_{m}))$, that reduces the distance between the kernel values of the samples in the data set and their Fourier feature approximation.
@@
where, $\mathcal{N}=\norm{\mathbf{z}_{\gamma/2}(\mathbf{x})}$, and the Fourier components constructed with parameter $\gamma/2$; it approximates the Gaussian kernel via $\innerproduct{\bar{\psi}(\mathbf{x})}{\bar{\psi}(\mathbf{y})}^{2}\approx e^{-\gamma{\norm{\mathbf{x}-\mathbf{y}}}}$ where |$\ket{\bar{\psi}(\mathbf{x})}\in\mathbb{R}^{d}$.
In this article, we propose the quantum random and quantum adaptive Fourier features, which map the space of the characteristics $\mathbf{x}\in\mathbb{R}^{D}$, to a complex Hilbert space |$\ket{\psi(\mathbf{x})}\in\mathbb{C}^{d}$, where the square of the modulus approximates the Gaussian kernel |$\absolutevalue{\innerproduct{\psi(\mathbf{x})}{\psi(\mathbf{y})}}^{2}\approx e^{-\gamma{\norm{\mathbf{x}-\mathbf{y}}}}$, see Sect. IV.1. For notational purposes, we use a bar to distinguish the quantum map based on RFF in the real domain |$\ket{\bar{\psi}(\mathbf{x})}$ from the quantum random and quantum adaptive Fourier features |$\ket{\psi(\mathbf{x})}$.
@@ III.3 Density Matrix Kernel Density Estimation
$=\frac{1}{NM_{\gamma}}\sum_{j=0}^{N-1}\absolutevalue{\innerproduct{\bar{\psi}(\mathbf{x}_{j})}{\bar{\psi}(\mathbf{x^{*}})}}^{2}$
$\approx\frac{1}{NM_{\gamma}}\sum_{j=0}^{N-1}e^{\gamma\norm{\mathbf{x}_{j}-\mathbf{x^{*}}}^{2}}=p_{\gamma}(\mathbf{x^{*}}).$ (6)
@@
$2^{8}\Big{(}\frac{\sqrt{2D\gamma}diam(\mathcal{M})}{3M_{\gamma}N\epsilon}\Big{)}\exp\Big(-\frac{d(3M_{\gamma}N\epsilon)^{2}}{4(D+2)}\Big{missing})$ (7)
Thus, the DMKDE algorithm provides a better approximation of the KDE method [14, 15] by increasing the number of Fourier features $d$ and the number of data samples $N$. This approximation has the advantage of reducing the computational complexity, since for $M$ samples, the DMKDE algorithm combined with RFF reduces the number of operations of the Parzen-Rosenblatt window from $O(NM)$ to $O(M+N)$, see Sect. IV.4, and its complexity does not grow exponentially with the size D of the data, in contrast to the fast multipole method for KDE [53, 54, 55].
@@ III.5 Amplitude encoding, unitaries and isometries
The application of an arbitrary unitary transformation $U\in\mathbb{C}^{d\times d}$ on $n=\lceil\log{(d)}\rceil$ ) ⌉ qubits is an important quantum subroutine, whose depth is of the order $O(d^{2})$. Unitary transformations are a subset of a more general transformation called isometries [56], which correspond to transformations from two Hilbert spaces that may have different dimensions, say $\mathbb{C}^{r}\rightarrow\mathbb{C}^{d}$. An isometry from $m=\log{(r)}$ ) to $n=\log{(d)}$ ) qubits can be regarded as a $\mathbb{C}^{d\times r}$ transformation built from the columns of a $\mathbb{C}^{d\times d}$ unitary matrix; amplitude encoding can also be seen as an isometry. Isometries are of particular interest because the circuit depth of isometries scales in the form $O(dr)$, which has lower complexity than arbitrary unitary transformations. Table 1 summarizes the number of qubits and circuit depth of amplitude encoding, unitary matrices, and isometries.
@@ IV.1 Quantum random Fourier features
To build the quantum features of the quantum density estimation method, we propose a novel embedding called quantum random Fourier features (QRFF), which correspond to RFF [28] in the complex domain. QRFF build a quantum feature map that approximates a Gaussian kernel by mapping classical data samples $\mathbf{x}\in\mathbb{R}^{D}$ to quantum state-like representations |$\ket{\psi(\mathbf{x})}\in\mathbb{C}^{d}$, such that, $\hat{k}(\mathbf{x},\mathbf{y})=\absolutevalue{\innerproduct{\psi(\mathbf{x})}{\psi(\mathbf{x})}{\psi(\mathbf{y})}}^{2}$ approximates $k(\mathbf{x},\mathbf{y})=e^{-\gamma{\norm{\mathbf{x}-\mathbf{y}}}^{2}}$. The QRFF is distinct from conventional RFF [28] by approximating the kernel via |$\absolutevalue{\innerproduct{\psi(\mathbf{x})}{\psi(\mathbf{x})}{\psi(\mathbf{y})}}^{2}$ instead of a dot product $\mathbf{z}_{\gamma}(\mathbf{x})^{T}\mathbf{z}_{\gamma}(\mathbf{y})$.
@@
$(\pi/\gamma)^{-(D/4)}\psi(\mathbf{x}-\mathbf{y})=$
@@
by writing the last integral as a sum, we have that, $E[e^{i\mathbf{p}\cdot\mathbf{x}}e^{-i\mathbf{p}\cdot\mathbf{y}}]\approx\innerproduct{\psi(\mathbf{x})}{\psi(\mathbf{x})}{\psi(\mathbf{y})}$ with |$\ket{\psi(\mathbf{x})}=\sqrt{1/d}\sum_{i=0}^{d-1}e^{i\mathbf{p}_{i}\cdot\mathbf{x}}\ket{i}$ is an unbiased estimate of $(\pi/\gamma)^{-(D/4)}\psi(\mathbf{x}-\mathbf{y})$, when $\{\mathbf{p}_{i}\}$ is sampled i.i.d. from $\mathcal{N}(\mathbf{0},\sqrt{\gamma}\mathbf{I}_{D})$, since the term that complements the complex exponential in Eq. 10 is a multivariate Gaussian with mean $\mathbf{0}$ and covariance matrix $\sqrt{\gamma}\mathbf{I}_{D}$. Therefore, we may built the Gaussian kernel by $k(\mathbf{x}-\mathbf{y})=e^{-\gamma{\norm{\mathbf{x}-\mathbf{y}}}^{2}}=(\pi/\gamma)^{-D/2}\psi(\mathbf{x}-\mathbf{y})^{*}\psi(\mathbf{x}-\mathbf{y})\approx\absolutevalue{\innerproduct{\psi(\mathbf{x})}{\psi(\mathbf{x})}{\psi(\mathbf{y})}}^{2}$, considering that, $\psi(\mathbf{x}-\mathbf{y})=(\gamma/\pi)^{-(D/4)}e^{-\gamma\norm{\mathbf{x}-\mathbf{y}}^{2}/2}$, for the Gaussian wavepacket in the position representation.
@@ IV.2 Quantum adaptive Fourier features
$\mathbf{w}_{j}^{(t+1)}=\mathbf{w}_{j}^{(t)}-\eta\nabla_{\mathbf{w}_{j}}\mathcal{L}_{\{\mathbf{w}_{k}\}},$ (12)
@@
To train the QAFF, we can use a siamese neural network to estimate |$\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}$ and to fit that value as close as possible to $e^{-\gamma{\norm{\bm{x}_{l}-\bm{x}_{m}}}^{2}}$, as in conventional AFF [32], see Fig. 2 for its neural architecture, however, in the Appendix. B, we present an explicit mathematical expression of the gradient $\nabla_{\mathbf{w}_{j}}\mathcal{L}_{\{\mathbf{w}_{k}\}}$, which avoids the use of the siamese neural network, also showing that the complexity of training QAFF in a classical computer is $O(TdD+Td^{2})$.
@@ IV.2.1 The kernel training data set to optimize the QAFF
The Gaussian kernel training data set: The main kernel training data set used in this article is the Gaussian data set. To approximate the Gaussian kernel $e^{-\gamma{\norm{\bm{x}_{l}-\bm{x}_{m}}}^{2}}$, whose original space is $\bm{x}\in\mathbb{R}^{D}$, we build a synthetic data set $\{\bm{x}_{l}\}_{1\cdots\mathcal{M}}$ sampled from a multivariate normal distribution $\mathcal{N}(\mathbf{0},\frac{1}{\sqrt{2\gamma}D}\mathbf{I}_{D})$. In addition, since the Gaussian kernel is shift-invariant, in this kernel training strategy we set one of the elements of each data pair equal to $\mathbf{0}$, so the new data pairs are $\{(\bm{x}_{l},\mathbf{0})\}_{1\cdots\mathcal{M}}$, and the loss function to optimize becomes $\mathcal{L}_{\{\mathbf{w}_{j}\}}=\frac{1}{\mathcal{M}}\sum_{l}(\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\mathbf{0})}}^{2}-e^{-\gamma\norm{\bm{x}_{l}}^{2}})^{2}$. The motivation of this kernel learning strategy is that the Gaussian data resulting from sampling $\mathcal{N}(\mathbf{0},\frac{1}{\sqrt{2\gamma}D}\mathbf{I}_{D})$ is equivalent to sampling from the kernel function $e^{-\gamma{\norm{\bm{x}}^{2}}}$, so there would be a greater number of kernel training points in the regions with larger values of the kernel function, leading to better approximation of the kernel in those regions. Furthermore, since $\gamma$ acts as a scaling factor in this kernel learning strategy, the learned weights do not depend on $\gamma$, and we can simplify the method by setting $\gamma=0.5$, nonetheless, when predicting the QAFF mapping on the training data set $\{\mathbf{x}_{j}\}_{1\cdots N}$ and testing sample $\mathbf{x^{*}}$, which are distinct from the kernel training data samples, it is necessary to choose the appropriate $\gamma$ value that approximates the desired kernel as in QRFF.

### 18 Mar 2024 v5

@@ I Introduction
In this article, we present a novel non-parametric quantum-classical density matrix density estimation strategy, called Q-DEMDE, suitable for current noisy quantum computers. It combines a proposed quantum algorithm for computing the expectation values of density matrices with a novel quantum representation of data called quantum Fourier features. The proposed method, unlike previous quantum density estimation methods [19, 20], builds probability distributions of data in the form of mixed quantum states in qubit-based quantum hardware. It is based on the density matrix kernel density estimation (DMKDE) algorithm [27] originally developed for classical computers; it fuses density matrices with random Fourier features (RFF) [28] to build a non-parametric density estimator. We show that the proposed Q-DEMDE method can be implemented in current IBM quantum computers.
@@ II Related Work
Density estimation is an important challenge in both classical and quantum machine learning. Both parametric and non-parametric methods for DE have been widely explored in classical computers [14, 15, 27, 16, 11, 12, 13]. Some early works on quantum density estimation build parametric models based on multivariate Gaussian distributions on quantum devices [17, 18]. On the non-parametric side, it is possible to combine the formalism of density matrices and kernels for density estimation as in the classical algorithms of kernel density matrices (KDM) [33] and DMKDE [27]; the latter combining density matrices with random Fourier features. Although DMKDE was originally developed for classical computers, it can be implemented in quantum computers with both pure-state and mixed-state approaches; the pure-state approach has been developed in qubit-based quantum computers [20], while the mixed-state version has been implemented in a high-dimensional quantum computer [19]. In contrast to these previous works, we present a QDE strategy built on mixed states for present-day qubit-based quantum devices and propose a quantum algorithm for estimating the expected value of a density matrix that extend the mixed-state DMKDE [19] to qubit-based quantum hardware.
@@ III.1 Kernel density estimation
$p_{\gamma}(\mathbf{x^{*}})=\frac{1}{NM_{\gamma}}\sum_{j=0}^{N-1}k(\mathbf{x}_{j},\mathbf{x^{*}})=\frac{1}{NM_{\gamma}}\sum_{j=0}^{N-1}e^{-\gamma\norm{\mathbf{x}_{j}-\mathbf{x^{*}}}^{2}},$ (1)
where $M_{\gamma}=(\pi/\gamma)^{\frac{D}{2}}$ is a normalization constant. This method scales with the size of the training data set and requires storing the whole data set in memory.
@@ III.2 Random and Adaptive Fourier features
by sampling i.i.d the weights $\{\mathbf{w}_{i}\}\in\mathbb{R}^{D}$ from $\mathcal{N}(\mathbf{0},\mathbf{I})$, and i.i.d. biases $\{b_{i}\}\in\mathbb{R}$ from $\text{Uniform}[0,2\pi]$, leading to $\mathbf{z}_{\gamma}(\mathbf{x})^{T}\mathbf{z}_{\gamma}(\mathbf{y})\approx e^{-\gamma{\norm{\mathbf{x}-\mathbf{y}}}^{2}}$. This method follows from Bochner theorem [53], which states that a shift-invariant kernel is the Fourier transform of a probability measure, e.g., the Gaussian kernel with parameter $\gamma$ is the Fourier transform of a normal distribution with variance $2\gamma$. In addition, the kernel approximation can be improved by means of adaptive Fourier features [32], which learns the optimal RFF weights and biases $\{(\mathbf{w}_{i},b_{i})\}$ by minimizing the loss function, $\mathcal{L}_{\{(\mathbf{w}_{i},b_{i})\}}=\sum_{l,m}(k(\mathbf{x}_{l},\mathbf{x}_{m})-\mathbf{z}_{\gamma}(\mathbf{x_{l}})^{T}\mathbf{z}_{\gamma}(\mathbf{x}_{m}))^{2}$, that reduces the distance between the kernel values of the samples in the data set and their Fourier feature approximation.
@@
where, $\mathcal{N}=\norm{\mathbf{z}_{\gamma/2}(\mathbf{x})}$, and the Fourier components constructed with parameter $\gamma/2$; it approximates the Gaussian kernel via $\langle\bar{\psi}(\mathbf{x})|\bar{\psi}(\mathbf{y})\rangle^{2}\approx e^{-\gamma{\norm{\mathbf{x}-\mathbf{y}}}^{2}}$ where |$\ket{\bar{\psi}(\mathbf{x})}\in\mathbb{R}^{d}$.
In this article, we propose the quantum random and quantum adaptive Fourier features, which map the space of the characteristics $\mathbf{x}\in\mathbb{R}^{D}$, to a complex Hilbert space |$\ket{\psi(\mathbf{x})}\in\mathbb{C}^{d}$, where the square of the modulus approximates the Gaussian kernel |$\absolutevalue{\langle{\psi}(\mathbf{x})|{\psi}(\mathbf{y})\rangle}^{2}\approx e^{-\gamma{\norm{\mathbf{x}-\mathbf{y}}}^{2}}$, see Sect. IV.1. For notational purposes, we use a bar to distinguish the quantum map based on RFF in the real domain |$\ket{\bar{\psi}(\mathbf{x})}$ from the quantum random and quantum adaptive Fourier features |$\ket{\psi(\mathbf{x})}$.
@@ III.3 Density matrix kernel density estimation
$=\frac{1}{NM_{\gamma}}\sum_{j=0}^{N-1}\absolutevalue{\langle\bar{\psi}(\mathbf{x}_{j})|\bar{\psi}(\mathbf{x^{*}})\rangle}^{2}$
$\approx\frac{1}{NM_{\gamma}}\sum_{j=0}^{N-1}e^{-\gamma\norm{\mathbf{x}_{j}-\mathbf{x^{*}}}^{2}}=p_{\gamma}(\mathbf{x^{*}}).$ (6)
@@
$2^{8}\Big{(}\frac{\sqrt{2D\gamma}diam(\mathcal{M})}{3M_{\gamma}\epsilon}\Big{)}\exp\Big(-\frac{d(3M_{\gamma}\epsilon)^{2}}{4(D+2)}\Big{missing})$ (7)
Thus, the DMKDE algorithm as an approximation of the KDE method [14, 15] can be improved by increasing the number of Fourier features $d$. This approximation has the advantage of reducing the computational complexity, since for $M$ samples, the DMKDE algorithm combined with RFF reduces the number of operations of the Parzen-Rosenblatt window from $O(NM)$ to $O(M+N)$, see Sect. IV.4, and its complexity does not grow exponentially with the size $D$ of the data, in contrast to the fast multipole method for KDE [54, 55, 56].
@@ III.5 Amplitude encoding, unitaries and isometries
The application of an arbitrary unitary transformation $U\in\mathbb{C}^{d\times d}$ on $n=\lceil\log{(d)}\rceil$ ) ⌉ qubits is an important quantum subroutine, whose depth is of the order $O(d^{2})$. Unitary transformations are a subset of a more general transformation called isometries [57], which correspond to transformations from two Hilbert spaces that may have different dimensions, say $\mathbb{C}^{r}\rightarrow\mathbb{C}^{d}$. An isometry from $m=\log{(r)}$ ) to $n=\log{(d)}$ ) qubits can be regarded as a $\mathbb{C}^{d\times r}$ transformation built from the first $r$ columns of a $\mathbb{C}^{d\times d}$ unitary matrix; amplitude encoding can also be seen as an isometry with $r=1$. Isometries are of particular interest because the circuit depth of isometries scales in the form $O(dr)$, which has lower complexity than arbitrary unitary transformations. Table 1 summarizes the number of qubits and circuit depth of amplitude encoding, unitary matrices, and isometries.
@@ IV.1 Quantum random Fourier features
To build the quantum features of the quantum density estimation method, we propose a novel embedding called quantum random Fourier features (QRFF), which correspond to RFF [28] in the complex domain. QRFF build a quantum feature map that approximates a Gaussian kernel by mapping classical data samples $\mathbf{x}\in\mathbb{R}^{D}$ to quantum state-like representations |$\ket{\psi(\mathbf{x})}\in\mathbb{C}^{d}$, such that, $\hat{k}(\mathbf{x},\mathbf{y})=\absolutevalue{\langle\psi(\mathbf{x})|\psi(\mathbf{y})\rangle}^{2}$ approximates $k(\mathbf{x},\mathbf{y})=e^{-\gamma{\norm{\mathbf{x}-\mathbf{y}}}^{2}}$. The QRFF is distinct from conventional RFF [28] by approximating the kernel via |$\absolutevalue{\langle\psi(\mathbf{x})|\psi(\mathbf{y})\rangle}^{2}$ instead of a dot product $\mathbf{z}_{\gamma}(\mathbf{x})^{T}\mathbf{z}_{\gamma}(\mathbf{y})$.
@@
$(\pi/\gamma)^{D/4}\psi(\mathbf{x}-\mathbf{y})=$
@@
by writing the last integral as a sum, we have that, $E[e^{i\mathbf{p}\cdot\mathbf{x}}e^{-i\mathbf{p}\cdot\mathbf{y}}]\approx\langle\psi(\mathbf{y})|\psi(\mathbf{x})\rangle$ with |$\ket{\psi(\mathbf{x})}=\sqrt{1/d}\sum_{j=0}^{d-1}e^{i\mathbf{p}_{j}\cdot\mathbf{x}}\ket{j}$ is an unbiased estimate of $(\pi/\gamma)^{D/4}\psi(\mathbf{x}-\mathbf{y})$, when $\{\mathbf{p}_{j}\}$ is sampled i.i.d. from $\mathcal{N}(\mathbf{0},\gamma\mathbf{I}_{D})$, since the term that complements the complex exponential in Eq. 10 is a multivariate Gaussian with mean $\mathbf{0}$ and covariance matrix $\gamma\mathbf{I}_{D}$. Therefore, we may built the Gaussian kernel by $k(\mathbf{x}-\mathbf{y})=e^{-\gamma{\norm{\mathbf{x}-\mathbf{y}}}^{2}}=(\pi/\gamma)^{D/2}\psi(\mathbf{x}-\mathbf{y})^{*}\psi(\mathbf{x}-\mathbf{y})\approx\absolutevalue{\langle\psi(\mathbf{x})|\psi(\mathbf{y})\rangle}^{2}$, considering that, $\psi(\mathbf{x}-\mathbf{y})=(\gamma/\pi)^{D/4}e^{-\gamma\norm{\mathbf{x}-\mathbf{y}}^{2}/2}$, for the Gaussian wavepacket in the position representation.
@@ IV.2 Quantum adaptive Fourier features
$\mathbf{w}_{j}^{(t+1)}=\mathbf{w}_{j}^{(t)}-\eta\nabla_{\mathbf{w}_{j}}\mathcal{L}_{\{\mathbf{w}_{k}\}}^{(t)},$ (12)
@@
To train the QAFF, we can use a siamese neural network to estimate |$\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}$ and to fit that value as close as possible to $e^{-\gamma{\norm{\bm{x}_{l}-\bm{x}_{m}}}^{2}}$, as in conventional AFF [32], see Fig. 2 for its neural architecture, however, in the Appendix. B, we present an explicit mathematical expression of the gradient $\nabla_{\mathbf{w}_{j}}\mathcal{L}_{\{\mathbf{w}_{k}\}}$, which avoids the use of the siamese neural network, also showing that the complexity of training QAFF in a classical computer is $O(TdD+Td^{2})$.
@@ IV.2.1 The kernel training data set to optimize the QAFF
The Gaussian kernel training data set: The main kernel training data set used in this article is the Gaussian data set. To approximate the Gaussian kernel $e^{-\gamma{\norm{\bm{x}_{l}-\bm{x}_{m}}}^{2}}$, whose original space is $\bm{x}\in\mathbb{R}^{D}$, we build a synthetic data set $\{\bm{x}_{l}\}_{1\cdots\mathcal{M}}$ sampled from a multivariate normal distribution $\mathcal{N}(\mathbf{0},\frac{1}{2\gamma D}\mathbf{I}_{D})$. In addition, since the Gaussian kernel is shift-invariant, in this kernel training strategy we set one of the elements of each data pair equal to $\mathbf{0}$, so the new data pairs are $\{(\bm{x}_{l},\mathbf{0})\}_{1\cdots\mathcal{M}}$, and the loss function to optimize becomes $\mathcal{L}_{\{\mathbf{w}_{j}\}}=\frac{1}{\mathcal{M}}\sum_{l}(\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\mathbf{0})\rangle}^{2}-e^{-\gamma\norm{\bm{x}_{l}}^{2}})^{2}$. The motivation of this kernel learning strategy is that the Gaussian data resulting from sampling $\mathcal{N}(\mathbf{0},\frac{1}{2\gamma D}\mathbf{I}_{D})$ is equivalent to sampling from the kernel function $e^{-\gamma{\norm{\bm{x}}^{2}}}$, so there would be a greater number of kernel training points in the regions with larger values of the kernel function, leading to better approximation of the kernel in those regions. Furthermore, since $\gamma$ acts as a scaling factor in this kernel learning strategy, the learned weights do not depend on $\gamma$, and we can simplify the method by setting $\gamma=0.5$, nonetheless, when predicting the QAFF mapping on the training data set $\{\mathbf{x}_{j}\}_{1\cdots N}$ and testing sample $\mathbf{x^{*}}$, which are distinct from the kernel training data samples, it is necessary to choose the appropriate $\gamma$ value that approximates the desired kernel as in QRFF.